# The effect of accretion on scalar superradiant instability

### v1

@@ Appendix A Some useful formulae
$\delta{M}_{\mathrm{s}}^{(011)}$ $=-\frac{3\left(484+9\pi^{2}\right)\alpha_{2}^{16}\text{$\mu$ }\tau_{\text{acc}}}{327680\alpha_{0}}\approx-5.24\times 10^{-3}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{16},$ (102a)
$\delta{M}_{\mathrm{s}}^{(022)}$ $=-\frac{\left(1024+49\pi^{2}\right)\alpha_{2}^{20}\text{$\mu$ }\tau_{\text{acc}}}{128566206720\alpha_{0}}\approx-1.17\times 10^{-8}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{20},$ (102b)
$\delta{M}_{\mathrm{s}}^{(033)}$ $=-\frac{11\alpha_{2}^{24}\text{$\mu$ }\tau_{\text{acc}}}{246230475079680\alpha_{0}}\approx-4.47\times 10^{-14}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{24},$ (102c)
$\delta{M}_{\mathrm{s}}^{(044)}$ $=-\frac{\left(16777216+2029052025\pi^{2}\right)\alpha_{2}^{28}\text{$\mu$ }\tau_{\text{acc}}}{2791851562500000000000000000\alpha_{0}}\approx-7.18\times 10^{-18}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{28},$ (102d)
$\delta{M}_{\mathrm{s}}^{(055)}$ $=-\frac{7\left(210453397504+480020337225\pi^{2}\right)\alpha_{2}^{32}\text{$\mu$ }\tau_{\text{acc}}}{34079635094274801587650648080384000\alpha_{0}}\approx-1.02\times 10^{-21}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{32},$ (102e)
$\delta{M}_{\mathrm{s}}^{(066)}$ $=-\frac{\left(53876069761024+91415073021129\pi^{2}\right)\alpha_{2}^{36}\text{$\mu$ }\tau_{\text{acc}}}{16320026746151167744397770322620658483200\alpha_{0}}\approx-5.86\times 10^{-26}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{36},$ (102f)
$\delta{M}_{\mathrm{s}}^{(077)}$ $=-\frac{6171\left(17592186044416+28821275417025\pi^{2}\right)\alpha_{2}^{40}\text{$\mu$ }\tau_{\text{acc}}}{961449231303759399228736679876164935336171929600\alpha_{0}}\approx-1.94\times 10^{-30}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{40}.$ (102g)
@@
$\alpha_{1}^{(022)}$$\approx\left\{-\frac{5\cdot 3^{12}}{2^{7}\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{2}^{7}\tau_{\mathrm{acc}}\mu}{{3}^{12}\cdot{5}}\left(M_{\mathrm{s0}}^{(022)}\mu e^{-\frac{{2}^{7}\left(\alpha_{1}^{(011)}\right)^{13}\tau_{\mathrm{acc}}\mu}{{3}^{11}\cdot{5}\cdot{13}}}\right)^{{13}/{3}}\right]\right\}^{{1}/{13}},$ (103a)
$\alpha_{1}^{(033)}$$\approx\left\{-\frac{{2}^{13}\cdot{3}^{3}\cdot{5}^{3}\cdot{7}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{3}^{25/3}\tau_{\mathrm{acc}}\mu}{{2}^{56/3}\cdot{5}^{3}\cdot{7}}\left(M_{\mathrm{s0}}^{(033)}\mu e^{-\frac{\left(\alpha_{1}^{(022)}\right)^{17}\tau_{\mathrm{acc}}\mu}{{2}^{13}\cdot{3}^{2}\cdot{5}^{3}\cdot{7}\cdot{17}}}\right)^{{17}/{3}}\right]\right\}^{{1}/{17}},$ (103b)
$\alpha_{1}^{(044)}$$\approx\left\{-\frac{{3}^{6}\cdot{5}^{15}\cdot{7}^{3}}{{2}^{11}\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{2}^{25}\cdot{3}\tau_{\mathrm{acc}}\mu}{{5}^{15}\cdot{7}^{3}}\left(M_{\mathrm{s0}}^{(044)}\mu e^{-\frac{{2}^{11}\left(\alpha_{1}^{(033)}\right)^{21}\tau_{\mathrm{acc}}\mu}{{3}^{6}\cdot{5}^{15}\cdot{7}^{4}}}\right)^{7}\right]\right\}^{{1}/{21}},$ (103c)
$\alpha_{1}^{(055)}$$\approx\left\{-\frac{{2}^{4}\cdot{3}^{23}\cdot{5}^{2}\cdot{7}^{3}\cdot{11}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{5}^{44/3}\tau_{\mathrm{acc}}\mu}{{2}^{4}\cdot{3}^{23}\cdot{7}^{3}\cdot{11}}\left(M_{\mathrm{s0}}^{(055)}\mu e^{-\frac{\left(\alpha_{1}^{(044)}\right)^{25}\tau_{\mathrm{acc}}\mu}{{2}^{4}\cdot{3}^{22}\cdot{5}^{4}\cdot{7}^{3}\cdot{11}}}\right)^{{25}/{3}}\right]\right\}^{{1}/{25}},$ (103d)
$\alpha_{1}^{(066)}$$\approx\left\{-\frac{{3}^{8}\cdot{5}^{3}\cdot{7}^{19}\cdot{11}^{3}\cdot{13}}{{2}^{14}\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{2}^{14}\cdot{3}^{34/3}\cdot{5}^{20/3}\tau_{\mathrm{acc}}\mu}{{7}^{19}\cdot{11}^{3}\cdot{13}}\left(M_{\mathrm{s0}}^{(066)}\mu e^{-\frac{{2}^{14}\left(\alpha_{1}^{(055)}\right)^{29}\tau_{\mathrm{acc}}\mu}{{3}^{7}\cdot{5}^{3}\cdot{7}^{19}\cdot{11}^{3}\cdot{13}\cdot{29}}}\right)^{{29}/{3}}\right]\right\}^{{1}/{29}},$ (103e)
$\alpha_{1}^{(077)}$$\approx\left\{-\frac{{2}^{40}\cdot{3}^{10}\cdot{5}^{3}\cdot{7}^{2}\cdot{11}^{3}\cdot{13}^{3}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{3}\cdot{7}^{20}\tau_{\mathrm{acc}}\mu}{{2}^{51}\cdot{5}^{3}\cdot{11}^{3}\cdot{13}^{3}}\left(M_{\mathrm{s0}}^{(077)}\mu e^{-\frac{\left(\alpha_{1}^{(066)}\right)^{33}\tau_{\mathrm{acc}}\mu}{{2}^{40}\cdot{3}^{10}\cdot{5}^{3}\cdot{7}^{2}\cdot{11}^{3}\cdot{13}^{3}}}\right)^{{33}/{3}}\right]\right\}^{{1}/{33}}$ (103f)
@@
$\alpha_{2}^{(011)}$ $=\left(\frac{5120\sqrt{6}}{9\pi^{2}\mu\tau_{\text{acc}}+484\mu\tau_{\text{acc}}}\right)^{1/14}\approx 1.24\times\left(\mu\tau_{\text{acc}}\right)^{-1/14}$ (104a)
$\alpha_{2}^{(022)}$ $=\left(\frac{2410616376\sqrt{6}}{49\pi^{2}\mu\tau_{\text{acc}}+1024\mu\tau_{\text{acc}}}\right)^{1/18}\approx 2.32\times\left(\mu\tau_{\text{acc}}\right)^{-1/18},$ (104b)
$\alpha_{2}^{(033)}$ $=\left(\frac{2564900782080\sqrt{6}}{11\mu\tau_{\text{acc}}}\right)^{1/22}\approx 3.42\times\left(\mu\tau_{\text{acc}}\right)^{-1/22},$ (104c)
$\alpha_{2}^{(044)}$ $=\left(\frac{18695434570312500000000000\sqrt{6}}{2029052025\pi^{2}\mu\tau_{\text{acc}}+16777216\mu\tau_{\text{acc}}}\right)^{1/26}\approx 3.90\times\left(\mu\tau_{\text{acc}}\right)^{-1/26},$ (104d)
$\alpha_{2}^{(055)}$ $=\left(\frac{159748289504413132442112412876800\sqrt{6}}{3360142360575\pi^{2}\mu\tau_{\text{acc}}+1473173782528\mu\tau_{\text{acc}}}\right)^{1/30}\approx 4.32\times\left(\mu\tau_{\text{acc}}\right)^{-1/30},$ (104e)
$\alpha_{2}^{(066)}$ $=\left(\frac{510000835817223992012430322581895577600\sqrt{\frac{2}{3}}}{274245219063387\pi^{2}\mu\tau_{\text{acc}}+161628209283072\mu\tau_{\text{acc}}}\right)^{1/34}\approx 4.80\times\left(\mu\tau_{\text{acc}}\right)^{-1/34},$ (104f)
$\alpha_{2}^{(077)}$ $=\left(\frac{858436813664070892168514892746575835121582080\sqrt{6}}{59285363532820425\pi^{2}\mu\tau_{\text{acc}}+36187126693363712\mu\tau_{\text{acc}}}\right)^{1/38}\approx 5.30\times\left(\mu\tau_{\text{acc}}\right)^{-1/38}.$ (104g)
@@
$\alpha_{2}^{(111)}$ $=\left(\frac{864}{5\mu\tau_{\text{acc}}}\right)^{1/11}\approx 1.60\times\left(\mu\tau_{\text{acc}}\right)^{-1/11}$ (105a)
$\alpha_{2}^{(022)}$ $=\left(\frac{7971615}{14\mu\tau_{\text{acc}}}\right)^{1/15}\approx 2.42\times\left(\mu\tau_{\text{acc}}\right)^{-1/15},$ (105b)
$\alpha_{2}^{(033)}$ $=\left(\frac{5734400000}{\mu\tau_{\text{acc}}}\right)^{1/19}\approx 3.26\times\left(\mu\tau_{\text{acc}}\right)^{-1/19},$ (105c)
$\alpha_{2}^{(044)}$ $=\left(\frac{190770721435546875}{1408\mu\tau_{\text{acc}}}\right)^{1/23}\approx 4.12\times\left(\mu\tau_{\text{acc}}\right)^{-1/23},$ (105d)
$\alpha_{2}^{(055)}$ $=\left(\frac{83543560665596539200}{13\mu\tau_{\text{acc}}}\right)^{1/27}\approx 4.97\times\left(\mu\tau_{\text{acc}}\right)^{-1/27},$ (105e)
$\alpha_{2}^{(066)}$ $=\left(\frac{35227172069319515266026045}{64\mu\tau_{\text{acc}}}\right)^{1/31}\approx 5.83\times\left(\mu\tau_{\text{acc}}\right)^{-1/31},$ (105f)
$\alpha_{2}^{(077)}$ $=\left(\frac{1339612089633916391635550208000}{17\mu\tau_{\text{acc}}}\right)^{1/35}\approx 6.69\times\left(\mu\tau_{\text{acc}}\right)^{-1/35}.$ (105g)

### v2

@@ Appendix A Some useful formulae
$\delta{M}_{\mathrm{s}}^{(011)}$ $=-\frac{3\left(484+9\pi^{2}\right)\alpha_{2}^{16}\text{$\mu$ }\tau_{\text{acc}}}{327680\alpha_{0}}\approx-5.24\times 10^{-3}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{16},$ (A.19a)
$\delta{M}_{\mathrm{s}}^{(022)}$ $=-\frac{\left(1024+49\pi^{2}\right)\alpha_{2}^{20}\text{$\mu$ }\tau_{\text{acc}}}{128566206720\alpha_{0}}\approx-1.17\times 10^{-8}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{20},$ (A.19b)
$\delta{M}_{\mathrm{s}}^{(033)}$ $=-\frac{11\alpha_{2}^{24}\text{$\mu$ }\tau_{\text{acc}}}{246230475079680\alpha_{0}}\approx-4.47\times 10^{-14}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{24},$ (A.19c)
$\delta{M}_{\mathrm{s}}^{(044)}$ $=-\frac{\left(16777216+2029052025\pi^{2}\right)\alpha_{2}^{28}\text{$\mu$ }\tau_{\text{acc}}}{2791851562500000000000000000\alpha_{0}}\approx-7.18\times 10^{-18}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{28},$ (A.19d)
$\delta{M}_{\mathrm{s}}^{(055)}$ $=-\frac{7\left(210453397504+480020337225\pi^{2}\right)\alpha_{2}^{32}\text{$\mu$ }\tau_{\text{acc}}}{34079635094274801587650648080384000\alpha_{0}}\approx-1.02\times 10^{-21}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{32},$ (A.19e)
$\begin{split}\delta{M}_{\mathrm{s}}^{(066)}&=-\frac{\left(53876069761024+91415073021129\pi^{2}\right)\alpha_{2}^{36}\text{$\mu$ }\tau_{\text{acc}}}{16320026746151167744397770322620658483200\alpha_{0}}\\ &\approx-5.86\times 10^{-26}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{36},\end{split}$ (A.19f)
$\begin{split}\delta{M}_{\mathrm{s}}^{(077)}&=-\frac{6171\left(17592186044416+28821275417025\pi^{2}\right)\alpha_{2}^{40}\text{$\mu$ }\tau_{\text{acc}}}{961449231303759399228736679876164935336171929600\alpha_{0}}\\ &\approx-1.94\times 10^{-30}\frac{\mu{\tau}_{\mathrm{acc}}}{\alpha_{0}}\alpha^{40}.\end{split}$ (A.19g)
@@
$\alpha_{1}^{(022)}$ $\approx\left\{-\frac{5\cdot 3^{12}}{2^{7}\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{2}^{7}\tau_{\mathrm{acc}}\mu}{{3}^{12}\cdot{5}}\left(M_{\mathrm{s0}}^{(022)}\mu e^{-\frac{{2}^{7}\left(\alpha_{1}^{(011)}\right)^{13}\tau_{\mathrm{acc}}\mu}{{3}^{11}\cdot{5}\cdot{13}}}\right)^{{13}/{3}}\right]\right\}^{{1}/{13}},$ (A.20a)
$\alpha_{1}^{(033)}$ $\approx\left\{-\frac{{2}^{13}\cdot{3}^{3}\cdot{5}^{3}\cdot{7}}{\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{3}^{25/3}\tau_{\mathrm{acc}}\mu}{{2}^{56/3}\cdot{5}^{3}\cdot{7}}\left(M_{\mathrm{s0}}^{(033)}\mu e^{-\frac{\left(\alpha_{1}^{(022)}\right)^{17}\tau_{\mathrm{acc}}\mu}{{2}^{13}\cdot{3}^{2}\cdot{5}^{3}\cdot{7}\cdot{17}}}\right)^{{17}/{3}}\right]\right\}^{{1}/{17}},$ (A.20b)
$\alpha_{1}^{(044)}$ $\approx\left\{-\frac{{3}^{6}\cdot{5}^{15}\cdot{7}^{3}}{{2}^{11}\tau_{\mathrm{acc}}\mu}W_{-1}\left[-\frac{{2}^{25}\cdot{3}\tau_{\mathrm{acc}}\mu}{{5}^{15}\cdot{7}^{3}}\left(M_{\mathrm{s0}}^{(044)}\mu e^{-\frac{{2}^{11}\left(\alpha_{1}^{(033)}\right)^{21}\tau_{\mathrm{acc}}\mu}{{3}^{6}\cdot{5}^{15}\cdot{7}^{4}}}\right)^{7}\right]\right\}^{{1}/{21}},$ (A.20c)
$\begin{split}\alpha_{1}^{(055)}&\approx\left(\frac{{2}^{4}\cdot{3}^{23}\cdot{5}^{2}\cdot{7}^{3}\cdot{11}}{\tau_{\mathrm{acc}}\mu}\right)^{{1}/{25}}\\ &\hskip 14.22636pt\times\left\{-W_{-1}\left[-\frac{{5}^{44/3}\tau_{\mathrm{acc}}\mu}{{2}^{4}\cdot{3}^{23}\cdot{7}^{3}\cdot{11}}\left(M_{\mathrm{s0}}^{(055)}\mu e^{-\frac{\left(\alpha_{1}^{(044)}\right)^{25}\tau_{\mathrm{acc}}\mu}{{2}^{4}\cdot{3}^{22}\cdot{5}^{4}\cdot{7}^{3}\cdot{11}}}\right)^{{25}/{3}}\right]\right\}^{{1}/{25}},\end{split}$ (A.20d)
$\begin{split}\alpha_{1}^{(066)}&\approx\left(\frac{{3}^{8}\cdot{5}^{3}\cdot{7}^{19}\cdot{11}^{3}\cdot{13}}{{2}^{14}\tau_{\mathrm{acc}}\mu}\right)^{{1}/{29}}\\ &\hskip 14.22636pt\times\left\{-W_{-1}\left[-\frac{{2}^{14}\cdot{3}^{34/3}\cdot{5}^{20/3}\tau_{\mathrm{acc}}\mu}{{7}^{19}\cdot{11}^{3}\cdot{13}}\left(M_{\mathrm{s0}}^{(066)}\mu e^{-\frac{{2}^{14}\left(\alpha_{1}^{(055)}\right)^{29}\tau_{\mathrm{acc}}\mu}{{3}^{7}\cdot{5}^{3}\cdot{7}^{19}\cdot{11}^{3}\cdot{13}\cdot{29}}}\right)^{{29}/{3}}\right]\right\}^{{1}/{29}},\end{split}$ (A.20e)
$\begin{split}\alpha_{1}^{(077)}&\approx\left(\frac{{2}^{40}\cdot{3}^{10}\cdot{5}^{3}\cdot{7}^{2}\cdot{11}^{3}\cdot{13}^{3}}{\tau_{\mathrm{acc}}\mu}\right)^{1/33}\\ &\hskip 14.22636pt\times\left\{-W_{-1}\left[-\frac{{3}\cdot{7}^{20}\tau_{\mathrm{acc}}\mu}{{2}^{51}\cdot{5}^{3}\cdot{11}^{3}\cdot{13}^{3}}\left(M_{\mathrm{s0}}^{(077)}\mu e^{-\frac{\left(\alpha_{1}^{(066)}\right)^{33}\tau_{\mathrm{acc}}\mu}{{2}^{40}\cdot{3}^{10}\cdot{5}^{3}\cdot{7}^{2}\cdot{11}^{3}\cdot{13}^{3}}}\right)^{{33}/{3}}\right]\right\}^{{1}/{33}}.\end{split}$ (A.20f)
@@
$\alpha_{2}^{(011)}$ $=\left(\frac{5120\sqrt{6}}{9\pi^{2}\mu\tau_{\text{acc}}+484\mu\tau_{\text{acc}}}\right)^{1/14}\approx 1.24\times\left(\mu\tau_{\text{acc}}\right)^{-1/14}$ (A.21a)
$\alpha_{2}^{(022)}$ $=\left(\frac{2410616376\sqrt{6}}{49\pi^{2}\mu\tau_{\text{acc}}+1024\mu\tau_{\text{acc}}}\right)^{1/18}\approx 2.32\times\left(\mu\tau_{\text{acc}}\right)^{-1/18},$ (A.21b)
$\alpha_{2}^{(033)}$ $=\left(\frac{2564900782080\sqrt{6}}{11\mu\tau_{\text{acc}}}\right)^{1/22}\approx 3.42\times\left(\mu\tau_{\text{acc}}\right)^{-1/22},$ (A.21c)
$\alpha_{2}^{(044)}$ $=\left(\frac{18695434570312500000000000\sqrt{6}}{2029052025\pi^{2}\mu\tau_{\text{acc}}+16777216\mu\tau_{\text{acc}}}\right)^{1/26}\approx 3.90\times\left(\mu\tau_{\text{acc}}\right)^{-1/26},$ (A.21d)
$\alpha_{2}^{(055)}$ $=\left(\frac{159748289504413132442112412876800\sqrt{6}}{3360142360575\pi^{2}\mu\tau_{\text{acc}}+1473173782528\mu\tau_{\text{acc}}}\right)^{1/30}\approx 4.32\times\left(\mu\tau_{\text{acc}}\right)^{-1/30},$ (A.21e)
$\begin{split}\alpha_{2}^{(066)}&=\left(\frac{510000835817223992012430322581895577600\sqrt{\frac{2}{3}}}{274245219063387\pi^{2}\mu\tau_{\text{acc}}+161628209283072\mu\tau_{\text{acc}}}\right)^{1/34}\\ &\approx 4.80\times\left(\mu\tau_{\text{acc}}\right)^{-1/34},\end{split}$ (A.21f)
$\begin{split}\alpha_{2}^{(077)}&=\left(\frac{858436813664070892168514892746575835121582080\sqrt{6}}{59285363532820425\pi^{2}\mu\tau_{\text{acc}}+36187126693363712\mu\tau_{\text{acc}}}\right)^{1/38}\\ &\approx 5.30\times\left(\mu\tau_{\text{acc}}\right)^{-1/38}.\end{split}$ (A.21g)
@@
$\alpha_{2}^{(111)}$ $=\left(\frac{864}{5\mu\tau_{\text{acc}}}\right)^{1/11}\approx 1.60\times\left(\mu\tau_{\text{acc}}\right)^{-1/11}$ (A.22a)
$\alpha_{2}^{(022)}$ $=\left(\frac{7971615}{14\mu\tau_{\text{acc}}}\right)^{1/15}\approx 2.42\times\left(\mu\tau_{\text{acc}}\right)^{-1/15},$ (A.22b)
$\alpha_{2}^{(033)}$ $=\left(\frac{5734400000}{\mu\tau_{\text{acc}}}\right)^{1/19}\approx 3.26\times\left(\mu\tau_{\text{acc}}\right)^{-1/19},$ (A.22c)
$\alpha_{2}^{(044)}$ $=\left(\frac{190770721435546875}{1408\mu\tau_{\text{acc}}}\right)^{1/23}\approx 4.12\times\left(\mu\tau_{\text{acc}}\right)^{-1/23},$ (A.22d)
$\alpha_{2}^{(055)}$ $=\left(\frac{83543560665596539200}{13\mu\tau_{\text{acc}}}\right)^{1/27}\approx 4.97\times\left(\mu\tau_{\text{acc}}\right)^{-1/27},$ (A.22e)
$\alpha_{2}^{(066)}$ $=\left(\frac{35227172069319515266026045}{64\mu\tau_{\text{acc}}}\right)^{1/31}\approx 5.83\times\left(\mu\tau_{\text{acc}}\right)^{-1/31},$ (A.22f)
$\alpha_{2}^{(077)}$ $=\left(\frac{1339612089633916391635550208000}{17\mu\tau_{\text{acc}}}\right)^{1/35}\approx 6.69\times\left(\mu\tau_{\text{acc}}\right)^{-1/35}.$ (A.22g)